OpenStudy (anonymous):

pLEASE PLEASE PLEASE CHECK TO SEE IF IM CORRECT 14X^2-6X+2 14(3)^2-6(3)+2. NOW iM STUCK i THINK THE ANSWER IS 110

OpenStudy (anonymous):

(fog) (3)

OpenStudy (anonymous):

On the first term, you first you have to do the power before the multiplication so: 14(9)-18+2

OpenStudy (amistre64):

what is f and g to begin with?

OpenStudy (amistre64):

without knowing what f and g is, there is no way to tell if youve done the question correctly

OpenStudy (anonymous):

f(x)=2x g(x)= 7x^2-3+1

OpenStudy (amistre64):

fog means that we need to insert g into the x parts of f

OpenStudy (amistre64):

f(x)=2x f(g) = 2g ; and since g(x)= 7x^2-3+1 fog = 2(7x^2-3+1) i think your missing an x in the g part

OpenStudy (anonymous):

yes 7x^2+3x+1 sorry

OpenStudy (amistre64):

lets see if it goes right there v fog = 2(7x^2-3x+1) and then when x=3 fog = 2(7(3)^2-3(3)+1) fog = 2(7(9)-9+1) fog = 2(63-8) fog = 2(55) = 110 yes

OpenStudy (anonymous):

so now we are at 14x^2-6x+2 right

OpenStudy (amistre64):

you dont "need" to work the 2 thru ... it can wait till afterwards if it makes life easier

OpenStudy (anonymous):

omg I got it right lol, please dont go now I have to do (gof) (x)

OpenStudy (amistre64):

is f and g the same rules as above?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

lets put f into g then g(x)=7x^2-3x+1 g(f)=7f^2-3f+1 and f=2x .... g(f)=7(2x)^2-3(2x)+1

OpenStudy (anonymous):

so now it would be 7(f(x))^2-3f(x)+1 7(2x)^2-3(2x)+1 7 (2x)(2x)-3(2x)+1 28x^2-6x+1

OpenStudy (amistre64):

i like it

OpenStudy (anonymous):

yeah next week is my final for this class and then I will have my aa

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

now I ahve to do (gof)(-4)

OpenStudy (anonymous):

amistre64 this is the one that I get lost on because of the (-4) 28x^2-6x+1 28(-4)^2-6(-4)+1 then I am lost would it be 448-24+1=425